# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v1

@@ Abstract
Selective retrieval improves retrieval-augmented generation (RAG) by reducing distractions from low-quality retrievals and improving efficiency. However, existing approaches under-utilize the inherent knowledge of large language models (LLMs), leading to suboptimal retrieval decisions and degraded generation performance. To bridge this gap, we propose Self-Routing RAG (SR-RAG), a novel framework that binds selective retrieval with knowledge verbalization. SR-RAG enables an LLM to dynamically decide between external retrieval and verbalizing its own parametric knowledge. To this end, we design a multi-task objective that jointly optimizes an LLM on knowledge source selection, knowledge verbalization, and response generation. We further introduce dynamic knowledge source inference via nearest neighbor search to improve the accuracy of knowledge source decision under domain shifts. Fine-tuning three LLMs with SR-RAG significantly improves both their response accuracy and inference latency. Compared to the strongest selective retrieval baseline, SR-RAG reduces retrievals by 29% while improving the performance by 5.1%. Data and code will be publicly released at https://github.com/xiaowu0162/self-routing-rag.
@@ 1 Introduction
However, a core question remains overlooked by current selective retrieval research:
@@
When retrieval is abstained, existing methods use a standard yet simplistic fallback: letting the LLM directly generate the response (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024; Wu et al., 2024a). This design restricts the LLM from explicitly articulating its parametric knowledge before generating a response. We argue that this ability of knowledge verbalization, while seemingly subtle, critically impacts the success of selective retrieval. First, verbalizing internal knowledge expands the LLM’s capacity to answer without retrieval. Prior work demonstrates the ability of LLMs to directly generate high quality knowledge (Yu et al., 2023) as well as intermediate reasoning paths (Wei et al., 2022; Allen-Zhu & Li, 2025) to benefit the system’s performance. This is especially valuable under complex queries, where naive retrieval methods return irrelevant results and compute-intensive retrieval may return noisy contexts. Second, knowledge verbalization enables more accurate selective retrieval decisions. Existing works train retrieval policies by comparing RAG with direct LLM response (Wang et al., 2023; Wu et al., 2024a) or resorting to likelihood preferences (He et al., 2021; Xu et al., 2024). By contrast, through explicit knowledge elicitation, knowledge verbalization helps characterize the LLM’s capabilities more precisely. Therefore, we argue that knowledge verbalization is a core component for selective retrieval to embrace.
@@
Figure 1: An overview of SR-RAG. Given a user query, the system first selects the most appropriate knowledge source by combining special token prediction with nearest neighbor search. Then, the knowledge is either retrieved from external an external source or self-verbalized by the LLM. Finally, the LLM forms the response based on the query and the knowledge. All the steps are streamlined as a single left-to-right generation pass.
@@
We propose Self-Routing RAG (SR-RAG), a selective retrieval framework that tightly integrates knowledge verbalization. By reformulating selective retrieval as a knowledge source selection problem, SR-RAG enables an LLM to self-route between retrieving external knowledge and verbalizing its own parametric knowledge, as illustrated in Figure 1. Recognizing the limitations of existing frameworks that fine-tune LLMs for selective retrieval via special token prediction (Asai et al., 2024; Wu et al., 2024a), SR-RAG introduces three key innovations. First, we generate diverse diverse verbalizations of the LLM’s internal knowledge to create more accurate training labels for knowledge source selection (Section 3.3). Second, we introduce a multi-task alignment objective that couples source selection, verbalization, and response generation. We further leverage self-supervised preference alignment over variants of verbalized knowledge to promote high-quality knowledge generation (Section 3.3). Finally, existing approaches suffer from poor source decision accuracy at inference time due to domain shifts and LLM ability shifts caused by fine-tuning. To bridge this gap, SR-RAG proposes dynamic knowledge source inference via nearest neighbor search, augmenting likelihood-based retrieval decisions with neighboring policy examples in the hidden representation space of the fine-tuned LLM (Section 3.4). Crucially, SR-RAG’s inference remains efficient, requiring only a single left-to-right generation pass.
@@
We evaluate SR-RAG by fine-tuning Llama-2-7B-Chat (Touvron et al., 2023), Phi-3.5-mini-instruct (Abdin et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) on a mixture of knowledge-intensive tasks to develop the LLMs’ ability to select knowledge sources and improve the quality of knowledge verbalization. Extensive experiments across four benchmarks demonstrate that SR-RAG greatly outperforms both always retrieving and the baseline selective retrieval approach. Compared to the latter, SR-RAG achieves 8.5%/2.1%/4.7% higher overall performance while performing 26%/40%/21% fewer retrievals across the three LLMs respectively (Section 5.2). Further analyses reveal that SR-RAG improves both the accuracy of selective retrieval decisions (Section 5.3) and the overall inference efficiency (Section 5.4). Finally, we conduct thorough ablation studies that confirm that the necessity of all three core components to SR-RAG’s strong performance (Section 5.5).
@@ 2 Related Work
To enhance the efficiency of RAG systems and avoid potentially harmful retrievals, several works have proposed to selectively skip retrieval augmentation, which we call selective retrieval following Xu et al. (2024) and Wu et al. (2024a). One popular approach is to assess whether retrieval augmentation increases the likelihood of the LLM generating the correct answer and distill this observation to a supervised decision model (He et al., 2021; Schick et al., 2023; Xu et al., 2024). Analogously, Wang et al. (2023) and Wu et al. (2024a) directly evaluate the correctness of the answers generated with and without retrieval to create the supervision signal. Recent works have also explored solely examining the question’s nature to judge the need for retrieval on question answering tasks (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024). To further incorporate LLM’s ability and confidence in the retrieval decision, Ding et al. (2024), Yao et al. (2024), and Moskvoretskii et al. (2025b) use the LLM’s uncertainty for selective retrieval. By contrast, this paper highlights the benefits of incorporating knowledge verbalization to make precise selective retrieval decisions and boosting the LLM’s performance when retrieval is skipped.
Adaptive RAG Inference Strategies
@@
A growing body of work has explored using LLMs to generate auxiliary knowledge for downstream tasks. Early studies demonstrated the potential of LLMs to produce relevant background knowledge in zero-shot settings for commonsense reasoning (Shwartz et al., 2020; Liu et al., 2022). More broadly, Yu et al. (2023) introduced a generate-then-read approach, treating LLMs as context generators to directly replace external retrieval for RAG. In addition, LLMs have been shown to generate effective intermediate reasoning steps for complex reasoning questions (Wei et al., 2022; Kojima et al., 2022). Building on these insights, this paper explores fully leveraging LLMs’ knowledge generation ability to benefit selective retrieval systems.
@@ 3.2 Self-Routing RAG: Overview
We propose Self-Routing RAG (SR-RAG), a generalization of selective retrieval that treats the LLM itself as a first-class knowledge source. As illustrated in Figure 1, given a query $q$, the LLM self-determines which knowledge source to use: either retrieving from an external source or verbalizing its parametric knowledge. The final response is then generated based on both the query and the knowledge collected from the selected source.
@@
    <EOQ> signals the end of the query and prompts the LLM for knowledge source selection [^1] This design aligns with Wu et al. (2024a) but diverges from Asai et al. (2024). We argue that this token is necessary for the LLM to allocate probability mass to the tokens for the knowledge sources..
@@
    <EOK> indicates the end of the knowledge and triggers the LLM for answer generation.
As shown in Figure 2, compared to traditional selective retrieval, SR-RAG enables the LLM to actively select the best knowledge source and seamlessly act as a knowledge source itself.
@@ 3.3 Self-Routing RAG: Training
To train the backbone LLM for SR-RAG, we propose a pipeline that mines self-supervision from widely available question answering or instruction following data, consisting of pairs of question $q$ and response $a$. Our pipeline only uses the LLM itself as the internal knowledge source $S_{i}$ and the external knowledge source $S_{e}$, without requiring additional human supervision or synthetic labels from stronger LLMs, demonstrating strong scalability.
@@
Then, each context $c_{j}$ is scored by the log-likelihood $l_{j}=p_{M}(a|q,c_{j})$ of generating the correct answer. Based on the ranking of $l_{j}$, we identify the preferred source $s\in\{S_{i},S_{e}\}$ as the one contributing to the majority of the top-$n$ ranked contexts. The resulting $(q,a,s,\{c_{j},l_{j}\})$ tuples are saved for model training. For convenience of later reference, we denote the contexts from $S_{i}$ and $S_{e}$ that lead to the highest and lowest likelihoods as $c_{i+},c_{i-},c_{e+},c_{e-}$, respectively. We present the formal data creation algorithm in Section B.1.
@@
$\mathcal{L}_{verb}^{DPO}=\begin{cases}-\log\sigma\Big{(}\beta\log\frac{p_{M}(c_{i+}|q)}{p_{ref}(c_{i+}|q)}-\beta\log\frac{p_{M}(c_{i-}|q)}{p_{ref}(c_{i-}|q)}\Big{)},&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$ (6)
$M$ and $ref$ are initialized with the LLM fine-tuned on $\mathcal{L}_{stage1}$, and only $M$ is updated.
Overall, this self-supervised pipeline effectively binds knowledge verbalization with the selective retrieval paradigm, enabling the LLM to learn accurate knowledge source preferences through performance-oriented labeling. Analogous to distilling complex “System 2” reasoning into fast “System 1” inference (Yu et al., 2024), the DPO objective leverages the computationally expensive high-quality knowledge to teach the LLM cost-efficient knowledge verbalization at inference time. Finally, SR-RAG naturally extends to more than two knowledge sources, which is useful in practice for distinguishing domain-specific corpora or retrieval methods with varying cost-quality trade-offs.
@@ 3.4 Self-Routing RAG: Inference
As shown in Figures 1 and 2, SR-RAG inference unfolds in a single left-to-right pass through three steps: source selection, knowledge collection, and answer generation.
@@
A common approach to selecting a knowledge source is to compare the likelihood $p_{M}(\text{{\textless s\textgreater}}|q)$ for each $s\in\mathcal{S}$, against a fixed threshold (Asai et al., 2024; Wu et al., 2024a). However, this approach fails to account for shifts in the LLM’s ability after fine-tuning and lacks fine-grained control over the decision boundary. To make source selection more robust, we propose a dynamic nearest neighbor-based source selection mechanism that builds a policy datastore based on rollout after the fine-tuning. Concretely, the fine-tuned LLM is first evaluated on a set of question-answer pairs [^2] Empirically, we reuse the training set so that no additional data or supervision is required.. For each pair, the probabilities of generating the answer $a$ conditioned on different knowledge sources are compared to decide the preferred one. We then build a policy datastore mapping each query to its preferred source, using the hidden representation at <EOQ> as the key. At inference, we retrieve $k$ nearest neighbors from the policy datastore and use their source labels to form a distribution over the sources $p_{D}(\text{{\textless s\textgreater}}|q)$. Finally, to select the best source $s\in\mathcal{S}$, we apply a threshold on the product:
@@
While tackling the challenges of source selection due to the LLM’s ability shift, this approach also exhibits better interpretability. Since the policy datastore consists of explicit source assignments, it can be audited, modified, and expanded by human experts to steer SR-RAG’s retrieval behavior in a fine-grained manner.
@@ 4.1 Implementation Details of SR-RAG
The main experiments are performed on two knowledge sources: the 2018 English Wikipedia (<Wiki>) as the external knowledge source, and the LLM itself (<Self>) as the internal knowledge source. We use the official Wikipedia embeddings released by Karpukhin et al. (2020) and retrieve in the granularity of 100-word chunks. GenRead (Yu et al., 2023) is used to verbalize diverse knowledge contexts. GenRead clusters zero-shot knowledge verbalizations in the same domain as in-context demonstrations to verbalize diverse knowledge. Each piece of verbalized knowledge is limited to a maximum of 150 tokens. From each knowledge source, we collect $n$ = 5 knowledge contexts.
@@
We fine-tune on a mixture of six short- and long-form knowledge-intensive datasets: Wizard of Wikipedia (Dinan et al., 2019), Natural Questions (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), OpenBookQA (Mihaylov et al., 2018), ARC-Easy (Bhakthavatsalam et al., 2021), and ASQA (Stelmakh et al., 2022). This mixture of 53,042 instances is a subset of the RAG instruction tuning data proven effective in Asai et al. (2024). After running the data construction algorithm discussed in Section 3.3 on three LLMs, 46.9% of the instances are labeled with <Self> and the rest are labeled with <Wiki> as the preferred knowledge source. Additional training details are provided in Section B.1.

### v2

@@ Abstract
Selective retrieval improves the accuracy and efficiency of retrieval-augmented generation (RAG) by reducing distractions from low-quality retrievals. However, existing approaches underutilize the inherent knowledge of large language models (LLMs), leading to suboptimal retrieval decisions and degraded generation performance. To bridge this gap, we propose Self-Routing RAG (SR-RAG), a novel framework that binds selective retrieval with knowledge verbalization. SR-RAG enables an LLM to dynamically decide whether to retrieve external knowledge or verbalize its own parametric knowledge. To this end, we design a multi-task objective that jointly optimizes an LLM for knowledge source selection, knowledge verbalization, and response generation. SR-RAG further incorporates a nearest neighbor search mechanism at inference time to improve the accuracy of knowledge source decisions under domain shifts. Fine-tuning three LLMs with SR-RAG significantly improves both their response accuracy and reduces the inference latency. Compared to the strongest selective retrieval baseline, SR-RAG reduces the number of retrievals by 29% while improving performance by 5.1%.
Self-Routing RAG: Binding Selective Retrieval with
Knowledge Verbalization
Di Wuthanks: Equal Contribution, Jia-Chen Gu11footnotemark: 1, Kai-Wei Chang, Nanyun Peng University of California, Los Angeles {diwu,kwchang,violetpeng}@cs.ucla.edu, gujc@ucla.edu
@@ 1 Introduction
However, a core question has been overlooked by current selective retrieval research:
@@
When retrieval is skipped, existing methods use a standard yet simplistic fallback: letting the LLM directly generate the response (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024; Wu et al., 2024a). This design prevents the LLM from explicitly articulating its parametric knowledge before generating a response. We argue that the ability to verbalize knowledge, while subtle, has a critical impact on the success of selective retrieval. First, verbalizing internal knowledge expands the LLM’s capacity to answer without retrieval. Prior work demonstrates the ability of LLMs to directly generate high-quality knowledge (Yu et al., 2023) as well as intermediate reasoning paths (Wei et al., 2022; Allen-Zhu and Li, 2025) to benefit the system’s performance. This is especially valuable for complex queries, where naive retrieval methods return irrelevant results and compute-intensive retrieval can return noisy contexts. Second, knowledge verbalization enables more accurate selective retrieval decisions. Existing works train retrieval policies by comparing RAG with direct LLM responses (Wang et al., 2023; Wu et al., 2024a) or resorting to likelihood-based preferences (He et al., 2021; Xu et al., 2024). By contrast, through explicit knowledge elicitation, knowledge verbalization helps characterize the LLM’s capabilities more precisely. Therefore, we argue that knowledge verbalization is a core component of selective retrieval to embrace.
@@
Figure 1: An overview of SR-RAG. Given a user query, the system first selects the most appropriate knowledge source by combining special token prediction with nearest neighbor search. Then, the knowledge is either retrieved from an external source or self-verbalized by the LLM. Finally, the LLM forms the response based on the query and the knowledge. All the steps are streamlined into a single left-to-right generation pass.
@@
We propose Self-Routing RAG (SR-RAG), a selective retrieval framework that tightly integrates knowledge verbalization. By reformulating selective retrieval as a knowledge source selection problem, SR-RAG enables an LLM to self-route between retrieving external knowledge and verbalizing its own parametric knowledge, as illustrated in Figure˜1. Recognizing the limitations of existing frameworks that fine-tune LLMs for selective retrieval via special token prediction (Asai et al., 2024; Wu et al., 2024a), SR-RAG introduces three key innovations. First, we generate diverse verbalizations of the LLM’s internal knowledge to create more accurate training labels for knowledge source selection (Section˜3.3). Second, we introduce a multi-task alignment objective that couples source selection, verbalization, and response generation. We further leverage self-supervised preference alignment over variants of verbalized knowledge to promote high-quality knowledge generation (Section˜3.3). Finally, existing approaches suffer from poor source decision accuracy at inference time due to domain shifts and LLM ability shifts caused by fine-tuning. To bridge this gap, SR-RAG proposes dynamic knowledge source inference via nearest neighbor search, augmenting likelihood-based retrieval decisions with neighboring policy examples in the hidden representation space of the fine-tuned LLM (Section˜3.4). Crucially, SR-RAG’s inference remains efficient, requiring only a single left-to-right generation pass.
@@
We evaluate SR-RAG by fine-tuning Llama-2-7B-Chat (Touvron et al., 2023), Phi-3.5-Mini-Instruct (Abdin et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024) on a mixture of knowledge-intensive tasks to develop the LLMs’ ability to select knowledge sources and improve the quality of knowledge verbalization. Extensive experiments across four benchmarks demonstrate that SR-RAG greatly outperforms both always retrieving and the baseline selective retrieval approach. Compared to the latter, SR-RAG achieves 8.5%/2.1%/4.7% higher overall performance while performing 26%/40%/21% fewer retrievals across the three LLMs respectively (Section˜5.2). Further analyses reveal that SR-RAG improves both the accuracy of selective retrieval decisions (Section˜5.3) and the overall inference efficiency (Section˜5.4). Finally, we conduct thorough ablation studies that confirm the necessity of all three core components to SR-RAG’s strong performance (Section˜5.5). Data and code will be publicly released at https://github.com/xiaowu0162/self-routing-rag.
@@ 2 Related Work
To enhance the efficiency of RAG systems and avoid potentially harmful retrievals, several works have proposed selectively skipping retrieval augmentation, which we call selective retrieval following Xu et al. (2024) and Wu et al. (2024a). One popular approach is to assess whether retrieval augmentation increases the likelihood of the LLM generating the correct answer and distill this observation into a supervised decision model (He et al., 2021; Schick et al., 2023; Xu et al., 2024). Analogously, Wang et al. (2023) and Wu et al. (2024a) directly evaluate the correctness of the answers generated with and without retrieval to create the supervision signal. Recent works have also explored solely examining the question’s nature to judge the need for retrieval on question answering tasks (Mallen et al., 2023; Jeong et al., 2024; Asai et al., 2024). To further incorporate the LLM’s confidence in the retrieval decision, Ding et al. (2024), Yao et al. (2024), and Moskvoretskii et al. (2025a) use uncertainty-based selective retrieval. By contrast, this paper highlights the benefits of incorporating knowledge verbalization to make precise selective retrieval decisions and to boost the performance when retrieval is skipped.
Adaptive RAG Inference
@@
A growing body of work has explored using LLMs to generate auxiliary knowledge. Early studies demonstrated the potential of LLMs to produce relevant background knowledge in zero-shot settings for commonsense reasoning (Shwartz et al., 2020; Liu et al., 2022). More broadly, Yu et al. (2023) introduced a generate-then-read approach, treating LLMs as context generators to replace external retrieval for RAG. In addition, LLMs have been shown to generate effective intermediate reasoning steps for complex reasoning questions (Wei et al., 2022; Kojima et al., 2022). Building on these insights, this paper explores fully leveraging LLMs’ knowledge generation ability to benefit selective retrieval systems.
@@ 3.2 Self-Routing RAG: Overview
We propose Self-Routing RAG (SR-RAG), a generalization of selective retrieval that treats the LLM itself as a first-class knowledge source. As illustrated in Figure˜1, given a query $q$, the LLM self-determines which knowledge source to use: either retrieving from an external source or verbalizing its parametric knowledge. The final response is then generated based on both the query and the knowledge collected from the selected source.
@@
    <EOQ> marks the end of the query and prompts the LLM for knowledge source selection [^1] This design aligns with Wu et al. (2024a) but diverges from Asai et al. (2024). We argue that this token is necessary for the LLM to allocate probability mass to the tokens for the knowledge sources..
@@
    <EOK> indicates the end of the knowledge and triggers the LLM to generate the answer.
As shown in Figure˜2, compared to traditional selective retrieval, SR-RAG enables the LLM to actively select the best knowledge source and seamlessly act as a knowledge source itself.
@@ 3.3 Self-Routing RAG: Training
To train the backbone LLM for SR-RAG, we propose a pipeline to mine self-supervision from widely available question answering or instruction-following data, consisting of pairs of question $q$ and response $a$. Our pipeline uses the LLM itself as the internal knowledge source $S_{i}$, alongside the external knowledge source $S_{e}$, without requiring additional human supervision or synthetic labels from stronger LLMs, demonstrating strong scalability.
@@
Then, each context $c_{j}$ is scored using the log-likelihood $l_{j}=p_{M}(a|q,c_{j})$ of generating the correct answer. Based on the ranking of $l_{j}$, we identify the preferred source $s\in\{S_{i},S_{e}\}$ as the one contributing to the majority of the top-$n$ ranked contexts. The resulting $(q,a,s,\{c_{j},l_{j}\})$ tuples are saved for model training. For convenience of later reference, we denote the contexts from $S_{i}$ and $S_{e}$ that lead to the highest and lowest likelihoods as $c_{i+},c_{i-},c_{e+},c_{e-}$, respectively. We present the formal data creation algorithm in Section˜B.1.
@@
$\mathcal{L}_{verb}^{DPO}=\begin{cases}-\log\sigma\Big(\beta\log\frac{p_{M}(c_{i+}|q)}{p_{ref}(c_{i+}|q)}-\beta\log\frac{p_{M}(c_{i-}|q)}{p_{ref}(c_{i-}|q)}\Big),&\text{if }s=S_{i},\\ 0,&\text{if }s=S_{e}.\end{cases}$
$M$ and refref are initialized with the LLM fine-tuned on $\mathcal{L}_{stage1}$, and only $M$ is updated.
Overall, this self-supervised pipeline effectively binds knowledge verbalization with the selective retrieval paradigm, enabling the LLM to learn accurate knowledge source preferences through performance-oriented labeling. Analogous to distilling complex “System 2” reasoning into fast “System 1” inference (Yu et al., 2024), the DPO objective leverages the computationally expensive high-quality knowledge to teach the LLM cost-efficient knowledge verbalization at inference time. Finally, SR-RAG naturally extends to more than two knowledge sources (further discussions and experiments in Appendix˜C), which is useful for distinguishing domain-specific corpora or retrieval methods with varying cost-quality trade-offs.
@@ 3.4 Self-Routing RAG: Inference
As shown in Figures˜1 and 2, SR-RAG inference unfolds in a single left-to-right pass through three steps: source selection, knowledge collection, and answer generation.
@@
A common approach to selecting a knowledge source is to compare the likelihood $p_{M}(\text{{\textless s\textgreater}}|q)$ for each $s\in\mathcal{S}$, against a fixed threshold (Asai et al., 2024; Wu et al., 2024a). However, this approach fails to account for shifts in the LLM’s ability after fine-tuning and lacks fine-grained control over the decision boundary. To make source selection more robust, we propose a dynamic nearest neighbor-based source selection mechanism that builds a policy datastore based on rollouts after the fine-tuning. Concretely, the fine-tuned LLM is first evaluated on a set of question-answer pairs [^2] Empirically, we reuse the training set so that no additional data or supervision is required.. For each pair, the probabilities of generating the answer $a$ conditioned on different knowledge sources are compared to decide the preferred source. We then build a policy datastore mapping each query to its preferred source, using the hidden representation at <EOQ> as the key. At inference, we retrieve $k$ nearest neighbors from the policy datastore and use their source labels to form a distribution over the sources $p_{D}(\text{{\textless s\textgreater}}|q)$. Finally, to select the best source $s\in\mathcal{S}$, we apply a threshold on the product:
@@
While tackling the challenges of source selection due to shifts in the LLM’s ability, this approach also exhibits better interpretability. Since the policy datastore consists of explicit source assignments, it can be audited, modified, and expanded by human experts to steer SR-RAG’s retrieval behavior in a fine-grained manner.
@@ 4.1 Implementation Details of SR-RAG
The main experiments are performed on two knowledge sources [^3] We present multi-source extensions of SR-RAG in Appendix C.: the 2018 English Wikipedia (<Wiki>) as the external knowledge source, and the LLM itself (<Self>) as the internal knowledge source. We use the official Wikipedia embeddings released by Karpukhin et al. (2020) and retrieve at a granularity of 100-word chunks. GenRead (Yu et al., 2023) is used to verbalize diverse knowledge contexts. GenRead clusters zero-shot knowledge verbalizations in as in-context demonstrations to verbalize diverse knowledge. We limit the verbalized knowledge chunks to a maximum of 150 tokens. From each knowledge source, we collect $n$ = 5 knowledge contexts.
@@
We fine-tune on a mixture of six short- and long-form knowledge-intensive datasets: Wizard of Wikipedia (Dinan et al., 2019), Natural Questions (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), OpenBookQA (Mihaylov et al., 2018), ARC-Easy (Bhakthavatsalam et al., 2021), and ASQA (Stelmakh et al., 2022). This mixture of 53,042 instances is a subset of the RAG instruction tuning data shown effective in Asai et al. (2024). After running the data construction algorithm, 46.9% of the instances are labeled with <Self> and the rest are labeled with <Wiki> as the preferred knowledge source, averaged across three LLMs. Full training details are provided in Section˜B.1.